OpenStudy (anonymous):

can some one help with this formula. I need to solve for b w=4/5(xb+m)

OpenStudy (phi):

\[ w = \frac{4}{5} \left( xb +m\right) \] the first step could be: multiply both sides by 5/4 what do you get ?

OpenStudy (anonymous):

5/4w=xb+m

OpenStudy (phi):

next add - m to both sides

OpenStudy (phi):

(but I would write it (5/4) w = xb+ m because 5/4w looks like \(\frac{5}{4w}\)

OpenStudy (phi):

or maybe 5w/4 = xb+m is clearer.

OpenStudy (phi):

next step is add - m to both sides 52/4 - m = xb + m - m

OpenStudy (phi):

next step is add - m to both sides 5w/4 - m = xb + m - m